OpenStudy (anonymous):

how do you approach the following problem? x/4= -x/2 -1

OpenStudy (anonymous):

Multiply both side by 4 to make it easier

OpenStudy (anonymous):

\[\frac{x}{4}=\frac{-x}{2}-1?\]

OpenStudy (anonymous):

the final solution should be -3/4

OpenStudy (anonymous):

thanks but i am still confused...can some one please show me the work done to solve this?

OpenStudy (anonymous):

my write is correct?

OpenStudy (anonymous):

x/4= - x/2 -1 fraction -x/2 is negative top and bottom... thnaks